What is a area of a circle 8.2
What is an answer
1
Expert's answer
2015-01-16T13:46:42-0500
The meaning of 8.2 is not indicated. If the length of circle`s radius is r=8.2, then the area will be πr2=π*8.22=211.24, where π is approximately equal to 3.14. If the length of circle`s diameter is d=8.2, then the area will be πr2=πd2/4=π*8.22/4=52.81, where π is approximately equal to 3.14.